# False vacuum decay and nucleation dynamics in neutral atom systems

### v1

@@ Abstract
False vacuum decay and nucleation offer the opportunity to study non-equilibrium dynamical phenomena in quantum many-body systems with confinement. Recent work has examined false vacuum decay in 1D ferromagnetic Ising spins and superfluids. In this paper, we study false vacuum nucleation dynamics in 1D antiferromagnetic neutral atom chains with Rydberg interactions, using both numerical simulations and analytic modeling. We apply a staggered local detuning field to generate the false and true vacuum states. Our efforts focus on two dynamical regimes: decay and annealing. In the first, we corroborate the phenomenological decay rate scaling and determine the associated parameter range for the decay process; in the second, we uncover and elucidate a procedure to anneal the false vacuum from the initial to the final system, with intermediate nucleation events. We further propose experimental protocols to prepare the required states and perform quenches on near-term neutral atom quantum simulators, examining the experimental feasibility of our proposed setup and parameter regime.
@@ I Introduction
False vacua are metastable states in quantum systems that lie above the true vacuum ground state and can decay by tunneling via nucleation [1, 2], thereby transforming under a first-order phase transition. They were examined in pioneering studies in quantum field theory [3, 4, 5, 1, 2, 6, 7], after earlier analysis in statistical physics [8, 9]. Metastable false vacua and their decay properties have also been explored in strongly-correlated quantum systems, including ultracold atoms [10, 11, 12] and magnetic materials [13, 14, 15]. In addition, more general metastable states in quasi-classical systems can also undergo phase transitions, including macromolecules such as polymers [16, 17] and proteins [18, 19], possibly influenced by similar effects.
@@
False vacuum metastability and decay are also important features of strongly-correlated quantum matter and materials. Experiments with these systems can be used to test field theoretic predictions and perhaps direct further theoretical investigations. Ultracold atoms are effective systems to study false vacua due to their controllability, and a number of experiments employing them have been proposed [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43] and implemented [10, 11, 12]. Notably, the authors of Ref. [12] recently performed an experiment that used ferromagnetic superfluids, formed from atomic Bose-Einstein condensates, and an adiabatic protocol to observe signatures of false vacuum decay driven by thermal and quantum fluctuations, realizing the first experimental test of the decay rate scaling. Quantum spin chains also offer many advantages for studying false vacuum phenomenology [13, 14, 15], particularly due to advances in the use of quantum quenches and adiabatic ramps [44, 45, 46]. A number of exegetic results have been obtained from these systems. The decay rate has been derived analytically in the perturbative limit [13], providing the spin analogue of the expression from quantum field theory. Numerical simulations with quantum quenches have recently exhibited and corroborated the exponential scaling [14]. More general false vacuum nucleation dynamics have been observed outside of the decay regime. Numerical investigations have examined linear ramps across first-order phase transitions and quantified the nucleation points for sequential domain sizes [47]. Real-time simulations have studied the collapse of false vacuum bubbles [48]. Scalar field theories have also been encoded in spin systems, with false vacua detected through simulated annealing [49].
@@ III.1.1 Quantum Field Theory
In quantum field theory, pioneering efforts to study false vacuum decay examined a scalar field Lagrangian $\mathcal{L}=\frac{1}{2}(\partial_{\mu}\phi)^{2}+V(\phi)$ [3, 4, 5, 1, 2, 6, 7, 86]. In the semiclassical limit where $\hbar$ is small, and thus the field is treated as a classical particle and the tunneling process is treated quantum mechanically, the decay rate, i.e. the tunneling probability per unit time, can be written as $\gamma=Ae^{-B}$, with $B=S_{E}[\phi_{b}]$ for $V(\phi_{FV})=0$, where the false vacuum is labeled $\phi_{FV}$ and the “bounce” solution $\phi_{b}$ is a stationary point of the Euclidean action $S_{E}$ [3, 5, 1, 2, 6, 86]. In the “thin-wall” limit where the critical bubble size is large compared to the bubble thickness, and thus the energy density difference $\mathcal{E}$ between the false and true vacua is small, the decay rate takes a simplified form since the rigorous upper and lower bounds on the decay exponent converge [87]. In the “thin-wall” limit, the form of the decay rate can be obtained in the semiclassical limit [3, 5, 1, 2, 6, 86] or, in the more general case, using effective action methods [4, 7]; in $1+1$ spacetime dimensions and in certain potentials, it has the form $\gamma\propto\mathcal{E}e^{-\lambda\mathcal{E}^{-1}}$ [5, 7], which becomes exponentially small for small $\mathcal{E}$.

### v2

@@ Abstract
Metastable states of quantum many-body systems with confinement offer a means to simulate false vacuum phenomenology, including non-equilibrium dynamical processes like decay by nucleation, in truncated limits. Recent work has examined the decay process in 1D ferromagnetic Ising spins and superfluids. In this paper, we study nucleation dynamics in 1D antiferromagnetic neutral atom chains with Rydberg interactions, using both numerical simulations and analytic modeling. We apply a staggered local detuning field to generate the metastable and ground states. Our efforts focus on two dynamical regimes: decay and annealing. In the first, we corroborate the phenomenological decay rate scaling and determine the associated parameter range for the decay process; in the second, we uncover and elucidate a procedure to anneal the metastable state from the initial to the final system, with intermediate nucleation events. We further propose experimental protocols to prepare the required states and perform quenches on near-term neutral atom quantum simulators, examining the experimental feasibility of our proposed setup and parameter regime.
@@ I Introduction
Metastable states in quantum systems lie above the true ground state and can decay by tunneling via nucleation [1, 2], thereby transforming under a first-order phase transition. They were examined in pioneering studies in quantum field theory as false vacua [3, 4, 5, 1, 2, 6, 7], after earlier analysis in statistical physics of similar phenomena [8, 9]. Metastable states and their decay properties have also been explored in strongly-correlated quantum systems, including ultracold atoms [10, 11, 12] and magnetic materials [13, 14, 15]. In addition, more general metastable states in quasi-classical systems can also undergo phase transitions, including macromolecules such as polymers [16, 17] and proteins [18, 19], possibly influenced by similar effects.
@@
Metastable states and their decay are also important features of strongly-correlated quantum matter and materials. Ultracold atoms are effective systems to study these states due to their controllability, and a number of experiments employing them have been proposed [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43] and implemented [10, 11, 12]. Notably, the authors of Ref. [12] recently performed an experiment that used ferromagnetic superfluids, formed from atomic Bose-Einstein condensates, and an adiabatic protocol to observe signatures of the decay process driven by thermal and quantum fluctuations, realizing the first experimental test of the decay rate scaling. Quantum spin chains also offer many advantages for studying metastability [13, 14, 15], particularly due to advances in the use of quantum quenches and adiabatic ramps [44, 45, 46]. A number of exegetic results have been obtained from these systems. The decay rate has been derived analytically in the perturbative limit [13], providing the spin analogue of the expression from quantum field theory. Numerical simulations with quantum quenches have recently exhibited and corroborated the exponential scaling [14]. More general nucleation dynamics have been observed outside of the decay regime. Numerical investigations have examined linear ramps across first-order phase transitions and quantified the nucleation points for sequential domain sizes [47]. Real-time simulations have studied the collapse of bubbles [48]. Scalar field theories have also been encoded in spin systems, with metastable states detected through simulated annealing [49].
The false and true vacuum states of quantum fields differ from the metastable and ground states of truncated quantum many-body systems, in that bosonic quantum fields include all field modes whereas finite-dimensional quantum systems have comparatively small Hilbert spaces. However, with these qualifications, we retain the former terms in the latter context due to the conceptual concordance between the two phenomena and the established terminology conventions in the literature (e.g. [14, 15, 48, 12]). Importantly, the decay process also occurs through nonperturbative bubble nucleation and the decay rate scaling exhibits a similar form. Experiments with these systems can investigate the reduction of field theoretical predictions to truncated models, and potentially motivate further theoretical investigations.
@@ III.1.1 Quantum Field Theory
In quantum field theory, pioneering efforts to study false vacuum decay examined a scalar field Lagrangian $\mathcal{L}=\frac{1}{2}(\partial_{\mu}\phi)^{2}-V(\phi)$ [3, 4, 5, 1, 2, 6, 7, 86]. In the semiclassical limit where $\hbar$ is small, and thus the field is treated as a classical particle and the tunneling process is treated quantum mechanically, the decay rate, i.e. the tunneling probability per unit time, can be written as $\gamma=Ae^{-B}$, with $B=S_{E}[\phi_{b}]$ for $V(\phi_{FV})=0$, where the false vacuum is labeled $\phi_{FV}$ and the “bounce” solution $\phi_{b}$ is a stationary point of the Euclidean action $S_{E}$ [3, 5, 1, 2, 6, 86]. In the “thin-wall” limit where the critical bubble size is large compared to the bubble thickness, and thus the energy density difference $\mathcal{E}$ between the false and true vacua is small, the decay rate takes a simplified form since the rigorous upper and lower bounds on the decay exponent converge [87]. In the “thin-wall” limit, the form of the decay rate can be obtained in the semiclassical limit [3, 5, 1, 2, 6, 86] or, in the more general case, using effective action methods [4, 7]; in $1+1$ spacetime dimensions and in certain potentials, it has the form $\gamma\propto\mathcal{E}e^{-\lambda\mathcal{E}^{-1}}$ [5, 7], which becomes exponentially small for small $\mathcal{E}$.